# On Linear Quadratic Potential Games

### v3

@@ Abstract
Our paper addresses characterizing conditions for a linear quadratic (LQ) game to be a potential game. The desired properties of potential games in finite action settings, such as convergence of learning dynamics to Nash equilibria, and the challenges of learning Nash equilibria in continuous state and action settings motivate us to characterize LQ potential games. Our first contribution is to show that the set of LQ games with full-state feedback that are potential games is very limited, essentially differing only slightly from an identical interest game. Given this finding, we restrict the class of LQ games to those with decoupled dynamics and decoupled state information structure. For this subclass, we show that the set of potential games strictly includes non-identical interest games and characterize conditions for the LQ games in this subclass to be potential. We further derive their corresponding potential function and prove the existence of a Nash equilibrium. Meanwhile, we highlight the challenges in the characterization and computation of Nash equilibrium for this class of potential LQ games.
@@
thanks: Corresponding author is Sara Hosseinirad
@@ 1 Introduction
Numerous emerging networked dynamical systems, such as smart grids and autonomous vehicles, rely on the decision-making of multiple interacting agents. Due to the increasing complexities of the dynamics and uncertainties of these systems, data-driven control has gained increasing attention. Reinforcement learning approaches like policy gradient methods have been widely applied in single-agent settings, and their convergence properties have been theoretically investigated [8]. However, multi-agent reinforcement learning has been less understood in both theory and practice.
@@
Most prior works on learning in LQ games have focused on policy gradient approaches. While the convergence of policy gradient and its variants in single-agent infinite-horizon LQ Gaussian control problems is well-established, the case of LQ games has proven more challenging. In particular, in infinite-horizon general-sum LQ game settings,  [16] and [17] argued that there are neither global nor local convergence guarantees to a Nash equilibrium for policy gradient. This negative result has motivated several follow-up studies on understanding the convergence of policy gradient in LQ games.
For finite-horizon general-sum LQ games, the global convergence of the natural policy gradient method to the Nash equilibrium is proven in [10]. However, this result relies on the assumption of sufficient noise for exploration. In [22], it was shown that the policy gradient algorithm has convergence guarantees when agents interact through the aggregated state and action. Other studies considered zero-sum [25], mean-field [5], or risk-sensitive [28] LQ games, proving the convergence of policy gradient variants in these settings.
In Markov games, of which LQ games are a subset, the challenges in proving convergence to Nash equilibria have motivated researchers to consider specific subclasses, particularly potential dynamic games. In static games, potential games are known for several desirable properties such as the existence of pure strategy Nash equilibria, the convergence of the best response dynamics [18], and learning dynamics to such equilibria [11]. These advances also have solid practical relevance in real-world dynamic games, including congestion games [21], resource allocation games [14], games arising in multi-vehicle autonomous driving [1, 13], and electricity markets [6, 20]. Consequently, a natural approach to extend the convergence guarantees of learning-based methods to dynamic games is to characterize the set of dynamic games that admit a potential function.
Recent literature in the multi-agent learning community has focused on conditions that ensure a dynamic game with a finite state and action spaces is potential. Studies on Markov games with finite states and actions [15, 26, 27] reveal that deriving conditions to characterize a potential Markov game is not straightforward. In particular, [15] demonstrated that the stage cost being potential is neither necessary nor sufficient for the dynamic game to be potential. Given these limitations, [26] focused on Markov games with decoupled dynamics and local policies where agents make decisions using only their local state. In this setting, they proved the convergence of a distributed learning algorithm to a Nash equilibrium for Markov potential games.
@@ 1.1 Contributions
  • For potential games within this subclass, we derive the potential function (Proposition 9) and show the existence of Nash equilibria (Proposition 10). Furthermore, we formalize the challenges associated with proving the uniqueness of the Nash equilibrium and its computation (Proposition 11 and Section 4.2).
The rest of the paper is organized as follows. In Section 2, we introduce the class of games under study and provide background information on potential games. In Section 3, we derive conditions for LQ potential games with coupled dynamic and full-state feedback and characterize the potential games in the case of decoupled dynamics and information structure. The properties of the LQ potential game, its corresponding structured optimal control problem, and the convergence of gradient descent for this control problem to a stationary point are provided in Section 4. We conclude and discuss open research directions in Section 5.
@@ 1.2 Notations
Herein, $\mathbb{R}$, $\mathbb{R}_{\geq 0}$ and $\mathbb{R}_{>0}$ refer to the sets of real numbers, non-negative, and positive real numbers, respectively. The summation of natural numbers $\{n^{p},\dots,n^{q}\}$ is denoted by $n^{p:q}=\sum_{j=p}^{q}n^{j}$. The set $[N]$ represents {1, …, N}, and $[N]\backslash\{i\}$ denotes all members of the set except $i$. The zero vector is denoted by $\mathbb{0}_{n}\in\mathbb{R}^{n}$, and the one vector by $\mathbb{1}_{n}\in\mathbb{R}^{n}$. We write $A=\text{blockdiag}\{A^{1},\dots,A^{N}\}$ to denote a block-diagonal matrix with $A^{1},\dots,A^{N}$ on its diagonal entries. For a vector $(\gamma^{1},\dots,\gamma^{N})\in\mathbb{R}^{n}$ where $n=n^{1}+\dots+n^{N}$, the vector $\gamma^{-i}\in\mathbb{R}^{\sum_{j\neq i}^{N}n^{j}}$ is defined as $\gamma^{-i}:=(\gamma^{1},\dots,\gamma^{i-1},\gamma^{i+1},\dots,\gamma^{N})$. The identity matrix of dimension $n$ is represented by $I_{n}$. Given a matrix $Q\in\mathbb{R}^{n\times n}$, $(Q)_{ij}$ is a sub-matrix of $Q$ that is formed by selecting entries from rows $(n^{1:i-1}+1)$ to $n^{1:i}$ and columns $(n^{1:j-1}+1)$ to $n^{1:j}$.
@@ 2.1 Linear quadratic games
Consider a non-cooperative general-sum LQ game where the dynamic is defined as
@@
$J^{i}_{u}(\gamma,x_{0}):=$ $\sum_{t=0}^{t_{f}-1}\sum_{j=1}^{N}\sum_{h=1}^{N}u_{t}^{j}(R^{i}_{t})_{jh}u^{h}_{t},$ (4)
with $\gamma^{i}=(\gamma^{i}_{0},\dots,\gamma^{i}_{t_{f}-1})\in\mathbb{R}^{t_{f}q^{i}}$ and $\gamma=(\gamma^{1},\dots,\gamma^{N})\in\mathbb{R}^{t_{f}q^{1:N}}$. The desired state at time $t$ is denoted by $d_{t}\in\mathbb{R}^{n}$. The parameters $Q_{t}^{i}\in\mathbb{R}^{n\times n}$ for $t\in\{0,\dots,t_{f}\}$ are positive semi-definite. For $t\in\{0,\dots,t_{f}-1\}$ and for $i,j,h\in[N]$, the cross-term action loss coefficients are $(R^{i}_{t})_{jh}=(R^{i}_{t})_{hj}\in\mathbb{R}$, and other action loss coefficients are $(R^{i}_{t})_{ll}\in\mathbb{R}_{\geq 0}$ for $l\in[N]\backslash\{i\}$ and $(R^{i}_{t})_{ii}\in\mathbb{R}_{>0}$.
@@
where $x_{t}^{i}\in\mathbb{R}^{n^{i}}$, $A^{i}\in\mathbb{R}^{n^{i}\times n^{i}}$, $b^{i}\in\mathbb{R}^{n^{i}}$ with $\sum_{i=1}^{N}n^{i}=n$. Note that the decoupled dynamics can be written as the dynamic introduced in (1) by defining the joint state as $x_{t}=(x_{t}^{1},\dots,x_{t}^{N})$, joint control as $u_{t}=(u_{t}^{1},\dots,u_{t}^{N})$, and $A=\text{blockdiag}\{A^{1},\dots,A^{N}\}\in\mathbb{R}^{n\times n}$, $B^{i}=(\mathbb{0}_{n^{1:i-1}},b^{i},\mathbb{0}_{n^{i+1:N}})\in\mathbb{R}^{n}$, for $i\in[N]$.
@@
Example 2: Let us consider a simple dynamic Cournot game, where agent $i$ is a firm that decides how much product $u^{i}_{t}\in\mathbb{R}$ to sell at each time step, given the amount of product stored $x^{i}_{t}\in\mathbb{R}_{\geq 0}$. A simple abstraction of the agent’s dynamic is
@@
The loss function parameters $Q_{t}^{i}=0$ for $t\in\{0,\dots,t_{f}-1\}$. In (3), the final stage loss parameter is $Q_{t_{f}}^{i}\in\mathbb{R}_{>0}$, and $d_{t_{f}}$ is the desired stored products at the final stage. In (4), the action loss coefficients $(R^{i}_{t})_{ji}=\alpha_{t}$ are the same for $i,j\in[N]$, and the rest of the action loss coefficients are $(R^{i}_{t})_{jh}=0$ for $j,h\in[N]\backslash\{i\}$.
@@ 3.1 Necessary and sufficient conditions for LQ potential games
3.2 Potential games with coupled dynamic
@@
Example 3. Here we consider an LQ game defined by equations (1) and (2) where $N=2$, $n=1$, and $d_{t}^{1}=d_{t}^{2}=0$, and the dynamic is $x_{t+1}=ax_{t}+b^{1}u^{1}_{t}+b^{2}u^{2}_{t}$ for $t\in\{0,\dots,t_{f}-1\}$. In loss function in (2), for open-loop information structure (i), $\gamma_{t}^{i}=u_{t}^{i}$, and for full-state feedback (ii), $\gamma_{t}^{i}=K_{t}^{i}$ where $u_{t}^{i}=-K_{t}^{i}x_{t}$, for $i\in\{1,2\}$.
@@
If $(R_{1}^{1})_{12}$ and $Q_{2}^{1}$ are respectively replaced by $(R_{1}^{2})_{12}$ and $Q_{2}^{2}$ in the equations above, the second derivatives of $J^{2}(\gamma)$ with respect to $u^{1}_{1}$ and $u^{2}_{\tau}$, for $\tau\in\{0,1\}$ are achieved. Condition in (10) is met for two pairs of $(t,\tau)\in\{(1,1),(1,0)\}$ if and only if $Q_{2}^{1}ab^{1}b^{2}=Q_{2}^{2}ab^{1}b^{2}$, and $Q_{2}^{1}b^{1}b^{2}+(R_{1}^{1})_{12}=Q_{2}^{2}b^{1}b^{2}+(R_{1}^{2})_{12}$. Since states are scalar, and dynamic parameters $b^{1}$ and $b^{2}$ are nonzero, these two conditions are met if and only if $Q_{2}^{1}=Q_{2}^{2}$ and $(R_{1}^{1})_{12}=(R_{1}^{2})_{12}$ that are equivalent to (12) for $t=1$.
@@
+$+\{(R^{1}_{1})_{11}u^{1}_{1}+(R^{1}_{1})_{12}u^{2}_{1}\}\frac{\partial u^{1}_{1}}{\partial K_{1}^{1}}\biggr{]}.$
@@
$\frac{\partial^{2}J^{1}(\gamma)}{\partial K_{0}^{2}\partial K_{1}^{1}}$ $=4\bigl{\{}Q_{2}^{1}ab^{1}b^{2}-[Q_{2}^{1}(b^{1})^{2}b^{2}+b^{2}(R_{1}^{1})_{11}]K_{1}^{1}$
$-[Q_{2}^{1}b^{1}(b^{2})^{2}+b^{2}(R_{1}^{1})_{12}]K_{1}^{2}\bigr{\}}\mathop{\mathbb{E}}_{x_{0}\sim\mathcal{D}}[x_{1}x_{0}].$
@@
+$+\{(R^{1}_{0})_{11}u^{1}_{0}+(R^{1}_{0})_{12}u^{2}_{0}\}\frac{\partial u^{1}_{0}}{\partial K_{0}^{1}}$
+$+\{(R^{1}_{1})_{11}u^{1}_{1}+(R^{1}_{1})_{12}u^{2}_{1}\}\frac{\partial u^{1}_{1}}{\partial K_{0}^{1}}$
+$+\{(R^{1}_{1})_{12}u^{1}_{1}+(R^{1}_{1})_{22}u^{2}_{1}\}\frac{\partial u^{2}_{1}}{\partial K_{0}^{1}}\biggr{]}.$
@@
$\frac{\partial^{2}J^{1}(\gamma)}{\partial K_{0}^{2}\partial K_{0}^{1}}$ $=2b^{1}b^{2}\big{[}Q_{1}^{1}+Q_{2}^{1}(a-b^{1}K_{1}^{1}-b^{2}K_{1}^{2})^{2}$
$+(R_{0}^{1})_{12}+(R_{1}^{1})_{11}(K_{1}^{1})^{2}+2(R_{1}^{1})_{12}K^{1}_{1}K^{2}_{1}$
$+(R_{1}^{1})_{22}(K_{1}^{2})^{2}\big{]}\mathop{\mathbb{E}}_{x_{0}\sim\mathcal{D}}[(x_{0})^{2}].$
@@
×$\times\mathop{\mathbb{E}}_{x_{0}\sim\mathcal{D}}[x_{h-1}x_{h-2}],$
@@
We expect that extending the result to higher dimensions will not in general relax the conditions derived for being a potential game. However, extending the proof beyond the scalar case is challenging, if not very tedious, due to reasoning about the set of equalities that arise. Let us consider $t_{f}=2$. Observe that for higher dimensions of states, a set of dynamic parameters might exist such that the condition in (10) is met for these pair of $(t,\tau)$ even though $Q_{2}^{1}\neq Q_{2}^{2}$. Such cases might happen when $B^{i}(B^{i})^{T}$ is not full rank, and consequently, the following condition derived through (10) for the pair of $(t,\tau)=(1,0)$ may not have a unique solution.
@@
The challenge of defining a class of potential games beyond identical interest on one hand and the desirable properties of a potential game on the other hand (see motivations discussed in the introduction) motivate us to confine the game setting to the decoupled dynamic defined in (5) and a decoupled state feedback information structure in (iii). In this setting, we identify LQ potential games that are not identical interest games.

### v4

@@ Abstract
Our paper addresses characterizing conditions for linear quadratic (LQ) games to be potential. The desired properties of potential games in finite action settings, such as convergence of learning dynamics to Nash equilibria and challenges of learning Nash equilibria in continuous state and action settings, motivate us to characterize LQ potential games. Our first contribution is to consider two-player LQ games with full-state feedback and scalar states and action space, and to analytically verify that the set of potential games within this example is limited, essentially differing only slightly from an identical interest game. Given this finding, we restrict the class of LQ games to those with decoupled dynamics and decoupled state linear feedback information structure. For this subclass, we show that the set of potential games strictly includes non-identical interest games and characterize conditions for LQ games in this subclass to be potential. We further derive their corresponding potential function and prove the existence of a Nash equilibrium. Meanwhile, we highlight the challenges in characterizations of the Nash equilibrium for this class of potential LQ games, theoretically and through simulation.
@@
thanks: This paper appears on Automatica
https://doi.org/10.1016/j.automatica.2025.112643
@@ 1 Introduction
Numerous emerging networked dynamic systems, such as smart grids and autonomous vehicles, rely on the decision-making of multiple interacting agents. Due to the increasing complexities of the dynamics and uncertainties of these systems, data-driven control has gained increasing attention. Reinforcement learning approaches like policy gradient methods have been widely applied in single-agent settings, and their convergence properties have been theoretically investigated [8]. However, multi-agent reinforcement learning has been less understood in both theory and practice.
@@
Most prior works on learning in LQ games have focused on policy gradient approaches. While the convergence of policy gradient and its variants in single-agent infinite-horizon LQ Gaussian control problems is well-established [8], the case of LQ games has proven more challenging. In particular, in infinite-horizon general-sum LQ game settings,  [16] and [17] argued that there are neither global nor local convergence guarantees to a Nash equilibrium for policy gradient. This negative result has motivated several follow-up studies on understanding the convergence of policy gradient in LQ games.
For finite-horizon general-sum LQ games, the global convergence of the natural policy gradient method to the Nash equilibrium is proven in [10]. However, this result relies on the assumption of sufficient noise for exploration. One study showed that the policy gradient algorithm has convergence guarantees when agents interact through the aggregated state and action [23]. Other studies considered zero-sum [26], mean-field [5], or risk-sensitive [29] LQ games, proving the convergence of policy gradient variants in these settings.
In Markov games, of which LQ games are a subset, the challenges in proving convergence to Nash equilibria have motivated researchers to consider specific subclasses, particularly potential dynamic games. In static games, potential games are known for several desirable properties such as the existence of pure strategy Nash equilibria, the convergence of the best response dynamics [18], and learning dynamics to such equilibria [11]. These advances also have solid practical relevance in real-world dynamic games. The LQ game is a well-known dynamic game setting for many real-world applications, including congestion games [22], resource allocation games [14], games arising in multi-vehicle autonomous driving [1, 13], and electricity markets [6, 20]. Consequently, a natural approach to extend the convergence guarantees of learning-based methods to dynamic games is to characterize the set of dynamic games that admit a potential function.
Recent literature in the multi-agent learning community has focused on conditions that ensure a dynamic game with finite state and action spaces is potential. Studies on Markov games with finite states and actions [15, 27, 28] reveal that deriving conditions to characterize a potential Markov game is not straightforward. In particular, [15] demonstrated that the stage cost being potential is neither necessary nor sufficient for the dynamic game to be potential. Given these limitations, [27] focused on Markov games with decoupled dynamics and local policies where agents make decisions using only their local state. In this setting, they proved the convergence of a distributed learning algorithm to a Nash equilibrium for Markov potential games. By characterizing LQ potential games as a subset of Markov potential games, we gain a fundamental understanding of a significant class of games in control theory.
@@ 1.1 Contributions
  • For potential games within this subclass, we derive the potential function (Proposition 11) and show the existence of Nash equilibria (Proposition 12). Furthermore, we formalize the challenges associated with proving the uniqueness of the Nash equilibrium and its computation theoretically (Proposition 13 and Section 4.2) and through simulation (Section 4.4).
The rest of the paper is organized as follows. In Section 2, we introduce the class of games under study and provide background information on potential games. In Section 3, we derive conditions for LQ potential games with coupled dynamics and full-state feedback and characterize the potential games in the case of decoupled dynamics and information structure. The properties of the LQ potential game, its corresponding structured optimal control problem, and the convergence of policy gradient for this control problem to a stationary point are provided in Section 4. We conclude and discuss open research directions in Section 5.
@@ 1.2 Notations
[…]
@@ 2.1 Linear quadratic games
Consider a non-cooperative general-sum LQ game where the dynamics are defined as
@@
$J^{i}_{u}(\gamma,x_{0}):=$ $\sum_{t=0}^{t_{f}-1}\sum_{j=1}^{N}\sum_{h=1}^{N}(u_{t}^{j})^{T}(R^{i}_{t})_{jh}u^{h}_{t},$ (4)
[…]
@@
where $x_{t}^{i}\in\mathbb{R}^{n^{i}}$, $A^{i}\in\mathbb{R}^{n^{i}\times n^{i}}$, $b^{i}\in\mathbb{R}^{n^{i}}$ with $\sum_{i=1}^{N}n^{i}=n$. Note that the decoupled dynamics can be written as the dynamics introduced in (22) by defining the joint state as $x_{t}=(x_{t}^{1},\dots,x_{t}^{N})$, joint control as $u_{t}=(u_{t}^{1},\dots,u_{t}^{N})$, and $A=\text{blockdiag}\{A^{1},\dots,A^{N}\}\in\mathbb{R}^{n\times n}$, $B^{i}=(\mathbb{0}_{n^{1:i-1}},b^{i},\mathbb{0}_{n^{i+1:N}})\in\mathbb{R}^{n}$, for $i\in[N]$.
@@
Example 2: Let us consider a simple dynamic Cournot game, where agent $i$ is a firm that decides how much product $u^{i}_{t}\in\mathbb{R}$ to sell at each time step, given the amount of product stored $x^{i}_{t}\in\mathbb{R}_{\geq 0}$. A simple abstraction of the agent’s dynamics is
@@
[…]
@@ 3.1 Necessary and sufficient conditions for LQ potential games
Note that each element of $\mathcal{J}(\gamma)$ is a second derivative of the loss function $J^{i}(\gamma)$, and the gradients of the loss functions can be more than second-order polynomials with respect to the decision variable, depending on the information structure. Therefore, investigating conditions for desirable properties in the Jacobian, such as diagonally strictly concave [21], can be as challenging as deriving LQ potential conditions. Next, we focus on investigating the potential condition in Lemma 4 for different game settings and information structures.
3.2 Potential games with coupled dynamics
@@
Example 3. Here we consider an LQ game defined by equations (22) and (2) where $N=2$, $n=1$, and $d_{t}^{1}=d_{t}^{2}=0$, and the dynamics is $x_{t+1}=ax_{t}+b^{1}u^{1}_{t}+b^{2}u^{2}_{t}$ for $t\in\{0,\dots,t_{f}-1\}$. In loss function in (2), for open-loop information structure (i), $\gamma_{t}^{i}=u_{t}^{i}$, and for full-state feedback (ii), $\gamma_{t}^{i}=K_{t}^{i}$ where $u_{t}^{i}=-K_{t}^{i}x_{t}$, for $i\in\{1,2\}$.
@@
If $(R_{1}^{1})_{12}$ and $Q_{2}^{1}$ are respectively replaced by $(R_{1}^{2})_{12}$ and $Q_{2}^{2}$ in the equations above, the second derivatives of $J^{2}(\gamma)$ with respect to $u^{1}_{1}$ and $u^{2}_{\tau}$, for $\tau\in\{0,1\}$ are achieved. Condition in (10) is met for two pairs of $(t,\tau)\in\{(1,1),(1,0)\}$ if and only if $Q_{2}^{1}ab^{1}b^{2}=Q_{2}^{2}ab^{1}b^{2}$, and $Q_{2}^{1}b^{1}b^{2}+(R_{1}^{1})_{12}=Q_{2}^{2}b^{1}b^{2}+(R_{1}^{2})_{12}$. Since states are scalar, and dynamics parameters $b^{1}$ and $b^{2}$ are nonzero, these two conditions are met if and only if $Q_{2}^{1}=Q_{2}^{2}$ and $(R_{1}^{1})_{12}=(R_{1}^{2})_{12}$ that are equivalent to (12) for $t=1$.
@@
$+\{(R^{1}_{1})_{11}u^{1}_{1}+(R^{1}_{1})_{12}u^{2}_{1}\}\frac{\partial u^{1}_{1}}{\partial K_{1}^{1}}\biggr].$
@@
$\frac{\partial^{2}J^{1}(\gamma)}{\partial K_{0}^{2}\partial K_{1}^{1}}$ $=4\big[Q_{2}^{1}b^{1}b^{2}(a-b^{1}K_{1}^{1}-b^{1}K_{1}^{2})$
$-b^{2}\{(R_{1}^{1})_{11}K_{1}^{1}+(R_{1}^{1})_{12}K_{1}^{2}\}\big]\mathop{\mathbb{E}}_{x_{0}\sim\mathcal{D}}[x_{1}x_{0}].$
@@
$+\{(R^{1}_{0})_{11}u^{1}_{0}+(R^{1}_{0})_{12}u^{2}_{0}\}\frac{\partial u^{1}_{0}}{\partial K_{0}^{1}}$
$+\{(R^{1}_{1})_{11}u^{1}_{1}+(R^{1}_{1})_{12}u^{2}_{1}\}\frac{\partial u^{1}_{1}}{\partial K_{0}^{1}}$
$+\{(R^{1}_{1})_{12}u^{1}_{1}+(R^{1}_{1})_{22}u^{2}_{1}\}\frac{\partial u^{2}_{1}}{\partial K_{0}^{1}}\biggr].$
@@
$\frac{\partial^{2}J^{1}(\gamma)}{\partial K_{0}^{2}\partial K_{0}^{1}}$ $=2\{b^{1}b^{2}\big[Q_{1}^{1}+Q_{2}^{1}(a-b^{1}K_{1}^{1}-b^{2}K_{1}^{2})^{2}$
$+(R_{1}^{1})_{11}(K_{1}^{1})^{2}+2(R_{1}^{1})_{12}K^{1}_{1}K^{2}_{1}$
$+(R_{1}^{1})_{22}(K_{1}^{2})^{2}\big]+(R_{0}^{1})_{12}\}\mathop{\mathbb{E}}_{x_{0}\sim\mathcal{D}}[(x_{0})^{2}].$
@@
$\times\mathop{\mathbb{E}}_{x_{0}\sim\mathcal{D}}[x_{h-1}x_{h-2}],$
@@
We expect that extending the result to higher dimensions will not, in general, relax the conditions derived for being a potential game. However, extending the proof beyond the scalar case is challenging, if not very tedious, due to reasoning about the set of equalities that arise. Let us consider $t_{f}=2$. Observe that for higher dimensions of states, a set of dynamics parameters might exist such that the condition in (10) is met for this pair of $(t,\tau)$ even though $Q_{2}^{1}\neq Q_{2}^{2}$. Such cases might happen when $B^{i}(B^{i})^{T}$ is not full rank, and consequently, the following condition derived through (10) for the pair of $(t,\tau)=(1,0)$ may not have a unique solution.
@@
Remark 8.
To investigate the effect of the noisy dynamics on the results of Proposition 5, we add a random noise $w_{t}$ drawn from the distribution $\mathcal{W}$ with zero mean to the game dynamics in equation (22). The loss function will be defined as follows.
$J^{i}(\gamma)=$ $\mathop{\mathbb{E}}_{x_{0}\sim\mathcal{D}}\bigl[\sum_{t=0}^{t_{f}-1}\mathop{\mathbb{E}}_{w_{t}\sim\mathcal{W}}c_{t}(x_{t+1},u_{t})\bigr],$ (15)
where the stage loss $c_{t}(x_{t+1},u_{t})$ is defined the same as the one for the game without noise. Considering the loss function defined above, the definition of Nash equilibrium, potential game, identical interest games for LQ games (Definitions 1, 2, and 3), the game pseudo-gradient, $\mathcal{G}(\gamma)$, and its Jacobian in equation (9) do not change. Hence, Lemma 4 is still valid for the game with noisy dynamics. Note that the derivative of $c_{t}(x_{t+1},u_{t})$ with respect to $x_{t}$ or $u_{t}^{j}$ is similar to the game without noise. For open-loop information structure, one can show that the derivatives of $x_{t+1}$ and $u_{t}^{i}$ with respect to the decision variable $u_{\tau}^{j}$ do not depend on $w_{t}$, which indicates that the results of Proposition 5, part (a) is the same for games with and without noise. For full-state linear feedback information structure, one can show that the derivatives of $x_{t+1}$ and $u_{t}^{i}$ with respect to the decision variable $K_{\tau}^{j}$ are zero for $t+1\geq\tau$ and for $t+1<\tau$ can be written as a product of a coefficient that do not depend on $w_{t}$ and $x_{\tau}$. This observation implies that by investigating conditions in Lemma 4, the potential conditions derived for the LQ game with noisy dynamics and full-state feedback information structure are the same as those for the game without noise in Proposition 5, part (b). Details are provided in Appendix A.
The challenge of defining a class of potential games beyond identical interest games, coupled with the desirable properties of potential games (as discussed in the introduction), motivates us to restrict the game setting. Given the properties of Markov potential games with decoupled dynamics and local policies in finite state and action spaces [27], we confine the LQ game setting, where the state and action spaces are infinite, to decoupled dynamics and the decoupled state linear feedback information structure. Moreover, restricting the $A$ and $B^{i}$ matrices in LQ games might lead to a potential game that satisfies the conditions in Remark 7 without being an identical interest game. As described in Section 2, LQ games with decoupled dynamics are a practically relevant subset of LQ games. On the other hand, the closed-loop Nash equilibrium of the general-sum LQ game, and consequently the LQ game with decoupled dynamics, has a linear full-state feedback structure [4]. To relax the conditions described in Proposition 5, an information structure sparser than full-state linear feedback, such as decoupled state linear feedback, should be investigated. Thus, we focus on LQ games with decoupled dynamics and a decoupled state linear feedback information structure, similar to the study [27] on Markov potential games with decoupled dynamics and localized policies.